OpenStudy (buckybarnes):

Help with paragraph proofs? Given: 1 and 2 are supplementary, and 2 and 3 are supplementary. Prove:1 is equal to 3 By the definition of supplementary angles, m1 + m2 = ______ (a) and m2 + m3 = ______ (b). Then m1 + m2 = m2 + m3 ______ (c). Subtract m2 from each side. You get m1 = _______ (d), or 1 _________ (e).

OpenStudy (buckybarnes):

My answers are A: 180 B: 180 C D: m3 E: 3

OpenStudy (buckybarnes):

@3mar

OpenStudy (3mar):

Peace be upon you.

OpenStudy (3mar):

Still did not get it?

OpenStudy (3mar):

I think it will be: C: substituting from previous equations

OpenStudy (3mar):

@triciaal

jhonyy9 (jhonyy9):

do you know what mean supplementary angles ?

OpenStudy (buckybarnes):

Yes

jhonyy9 (jhonyy9):

write me please here

jhonyy9 (jhonyy9):

1 and 2 are supplementary => what about them ?

jhonyy9 (jhonyy9):

=> 1 +2 = 180 degrees yes ?

OpenStudy (buckybarnes):

Yes

jhonyy9 (jhonyy9):

and given again that 2 + 3 = 180 degrees too yes ?

OpenStudy (buckybarnes):

Yes

jhonyy9 (jhonyy9):

so 1 +2=180 2 +3 = 180 ------------ => 1+2 = 2+3 subtract 2 from both sides and will get 1 = 3 hope helped

OpenStudy (buckybarnes):

I already knew that part. Thanks

jhonyy9 (jhonyy9):

ok. hope you understand it now good luck bye bye

OpenStudy (3mar):

She just want to know what is in the blank C, jhonyy9.

jhonyy9 (jhonyy9):

m1 = m3

jhonyy9 (jhonyy9):

after you subtract from both sides m2 you get this m1=m3

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (buckybarnes):

I understand that part too but that wasn't what I was asking for thank you for helping me though.